What do you mean by conservative forces? An elevator weighing 400 kg is to be lifted up at a constant velocity of 3 cm s–1, what would be the minimum power of the motor to be used?
1
Expert's answer
2013-07-12T09:34:11-0400
A conservative force is a force with the property that the work done in moving a body between two points is independent of the path taken. Equivalently, if a body travels in a closed loop, the net work done (the sum of the force acting along the path multiplied by the distance travelled) by a conservative force is zero.


To find power, you have to multiply applied force (F=mg) on velocity (v=0.03 m/s),
N=F*v = mg*v = 400*9.8*0.03 m/s = 117.6 Wt